# Reservoir-Computing Model for Mapping and Forecasting Neuronal Interactions from Electrophysiological Data

### 23 Jan 2024 v2

@@ 4.1 The paradigm
We consider a multi-site measurement of electrophysiological signals from a neuronal culture, such as 2D microelectrode array (MEA). We seek to represent the tested culture as a network where each node corresponds to one measurement electrode. Each electrode samples the electrophysiological signals from the neuron ensemble (consisting of a few neurons) found in its vicinity (Fig. 0(a)). Therefore, each node has to represent a complex neuronal circuit whose dynamics by itself is driven by numerous interacting neurons. We hence define the domain of the measurement as the macroscopic domain, which is described by the network in question; whereas the neuronal structure sampled by each node will be referred as the microscopic domain (or later as the reservoir domain). The data unit which is contained in each of these nodes is a sample of the electrophysiological signals expressed in instantaneous spike-rate measured in a specified time window. By “data unit” we refer to a set of data sampled at the network nodes in a definite time window, which contains information on the status of the network, with a memory on the previous time steps, and the ability to predict the next step accordingly. The time window is determined by a characteristic information flow rate of the network. This unit of time is dependent on many properties of the network such as neuron density in the culture, age of the culture and other [26], and it characterizes the signal integration time of each node.
@@
Let us represent the macro-domain state of the network at each time step $n=1,2,3...$ with a vector $\mathbf{y}[n]$, where each component of the vector describes the state of a single node, i.e., $\mathbf{y}[n]$ is the signal representation of each electrode at time $n$. Our aim is to seek the time propagation operator (or a synaptic transmission function) $\hat{\mathcal{F}}$ (Eq. (8)) transforming state $\mathbf{y}[n]$ to $\mathbf{y}[n+1]$. The operator $\hat{\mathcal{F}}$, which is likely to be non-linear due to the nature of neuronal networks, should describe as closely as possible the experimental observation in the electrophysiological measurements, i.e., we aim to fit a model to an observation which can mimic or predict the spatio-temporal patterns of the neuronal activity in the culture under test.
@@
We then consider the fact that each node in the macro-domain network represents a complex neuronal signal-processing-unit. It arises from the fact that typically every measurement site is surrounded by neurons which may be as many as a dozen. The morphology and functionality of each of these micro-circuits embedded in each node of the macro-domain network cannot be easily obtained from the electrophysiological measurements of standard recording systems. Also modeling of such neuronal structures is not an easy task and has been studied for decades, with numerous models for different scales of dimensions and time [4, 5].
@@
Hence our approach is to represent each measurement node (electrode) as a gate to a particular neuronal circuit (reservoir), where the signals measured at each node are an outcome of a complex operation involving each circuit and the whole network. We therefore propose the artificial neural network (ANN) structure depicted in Fig. 0(c) This structure represents a simplification of the neuronal dynamics, where each of the neuronal circuits is a black box, whose morphology and functionality are not known but assumed to be reasonably random.
@@
As seen in Fig. 0(c) and 0(d), we assume that the signal sampled at each node is an input to and an output from a higher dimensional domain with a specific connectivity and functionality. At the input of the ANN, the signal at each node is transformed to a corresponding reservoir-state (in the reservoir domain) by a set of uncoupled weighted connections (Input layer). Each micro-reservoir, associated with a node in the macro-domain, represents a micro-neuronal circuit embedded at each of the measurement sites, and has inner interconnections which represent the connectivity of the micro-circuits (Reservoir Layer). Each such circuit performs a nonlinear transformation, creating an updated reservoir state, which on one side is stored as a memory to be integrated to the next time steps, and on the other side is used to form the next state of the macro-domain network by weighting and coupling all the micro-reservoir states (Output Layer); then the whole process repeats cyclically. This kind of recurrent network is known as Reservoir Computing Network (RCN) and has been widely studied.
@@ 4.2.1 Domains and Dimensions
As mentioned above, the model distinguishes between two domains: The macro-domain which refers to the experimental observations, represented by the corresponding network; and the micro- (or reservoir) domain which refers to the neuronal units embedded in each of the macro network nodes, with no experimental data. We denote by $N_{ch}$ the dimension of the macro-network which in fact represents the number of nodes in the network, where each node is directly associated with an electrode (or a channel) in the experimental measurement. $N_{res}$ is the dimension of the reservoir.
Assuming that the neurons are uniformly distributed in the culture, we appoint a fixed number of connections between each node and the corresponding micro-circuit, such that for each node of the network there is one micro-reservoir (see Fig. 0(c)):
$N_{res}=mN_{ch}$ (10)
where $m$ is an integer number. It follows that each $m$ components in the vector space of the reservoir domain correspond to one node in the macro domain. In fact, we may associate $m$ with a relative size of each micro-circuit.
@@ 4.2.2 Input Layer
The input layer refers to the stage between the macro domain and the reservoir one. Here we assume that the data at each of the nodes is a linear transformation of the corresponding input state to the reservoir, such that each component in the macro-domain transforms directly to corresponding $m$ inputs of $N_{res}$ components in the reservoir domain, and refer to a single micro-circuit. This is done with the linear transformation described by Eq. (2a), where the state $\mathbf{y}\in\mathbb{R}^{N_{ch}\times 1}$ is transformed to the input reservoir state $\mathbf{x}_{in}\in\mathbb{R}^{N_{res}}$ through an input matrix $\mathcal{W}_{in}\in\mathbb{R}^{N_{res}\times N_{ch}}$. Since $\mathcal{W}_{in}$ maps each node to a corresponding micro-reservoir, it is represented by the following matrix:
${\mathcal{W}_{in}=\begin{bmatrix}\left(\mathbf{w}_{in}^{(1)}\right)\par&0&0&\cdots&0\\ 0&\left(\mathbf{w}_{in}^{(2)}\right)&0&\cdots&0\\ 0&0&\left(\mathbf{w}_{in}^{(3)}\right)&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\\ 0&0&0&&\left(\mathbf{w}_{in}^{{(N_{ch})}}\right)\end{bmatrix}}$ (11)
@@ 4.2.3 Reservoir Layer
The reservoir layer contains $N_{ch}$ independent micro-circuits with $m$ nodes each (total $N_{res}=m\times N_{ch}$ nodes). Each such circuit models the neuronal circuit around each electrode. This layer has two main functionalities: 1. nonlinear time-operator. 2. Reservoir state integrator. The dynamics of the reservoir layer is recurrent, where each time step $n$ the reservoir state $\mathbf{x}[n-1]$ is nonlinearly mapped into a new reservoir state $\mathbf{x}[n]$ (Eq. (2b)). This discrete differential relation provides cumulative data at each time step and carries the temporal memory on the activity of the network. Note that at time step $n$, the reservoir state $\mathbf{x}[n-1]$ undergoes an inner transformation through the reservoir matrix $\mathcal{W}_{res}\in\mathbb{R}^{N_{res}\times N_{res}}$, then integrated with the reservoir input state $\mathbf{x}_{in}$, followed by a nonlinear operation $\mathbf{f}_{NL}$. The operation of $\mathcal{W}_{res}$ is linear and describes the inner data propagation between the nodes of the reservoir. The synaptic morphology of each micro-reservoir is random and since they are independent, there are no coupling elements between them at this stage of the ANN. In addition, we assume that $\mathcal{W}_{res}$ preserves the energy of the state, hence it is represented by a block-diagonal and orthogonal matrix as the following:
@@
Further we note in Eq. (2b) the memory parameter $10<\alpha<$ 1, which factorizes the orthogonal matrix $\mathcal{W}_{res}$. This represent the memory strength, such that the effect of the reservoir state $\mathbf{x}[n]$ on the state $\mathbf{x}[n+k]$ will scale as $\alpha^{k}$. $\alpha=0$ indicates that the system is memoryless and the current state at time step $n$ depends only on the input.
The nonlinear function $\mathbf{f}_{NL}$ in Eq. (2b) accounts for the impact of network saturation and plasticity. Specifically, when subjected to a persistent input, the reservoir tends to saturate, while in the absence of input, it retains an excited state due to the influence of preceding signals (plasticity). Note that the length of the plasticity effect depends on the parameter $\alpha$. In this work we tried to implement different types of nonlinear functions with such effects, such as tanh and sigmoid, however the performance of the model showed a weak dependence on the choice of the function. For the results presented in this paper we considered:
@@
where $\Theta(\mathbf{\xi})$ is the 00 / $1$ step function at $\mathbf{\xi}=0$ used to limit the emerging values from the reservoir layer to non-negative only. Further we note that in Eq. (2b) the nonlinear function is $\mathbf{f}_{NL}(\hat{\mathcal{S}}\mathbf{\xi})$. The factor $\hat{\mathcal{S}}$ gives each reservoir neuron a different strength of the nonlinear response. This operation does not couple the different neurons and hence $\hat{\mathcal{S}}\in\mathbb{R}^{N_{res}\times N_{res}}$ is represented by a diagonal matrix containing normally-distributed synaptic strengths on its diagonal.
@@ 4.2.4 Output Layer
The output layer transforms the reservoir state back to the macroscopic domain. Here we assume a fully connected layer, such that all the $N_{res}$ reservoir nodes are weighted and connected to $N_{ch}$ nodes of the macroscopic network. This layer practically expresses the synaptic connectivity between the different nodes of the network. It is assumed that this transformation is purely linear, taking in consideration that the overall nonlinearity of the model is dominated by the reservoir layer. The relation of the output layer is given by Eq. (2c).
Unlike $\mathcal{W}_{in}$, $\mathcal{W}_{res}$ and $\hat{\mathcal{S}}$, which are matrices with random and constrained weights, $\mathcal{W}_{out}$ has no constraints on the values of its weights, rather it is the layer which is trained with Lasso regression (Eq. (2.1)).
@@ 4.3 Data
The model operates on the principles of a rate-coded spiking neural network. For model training, we generated tailored datasets that govern episodes of network bursting (NB) activity, representing them as multidimensional sequences of ISR data. The data in this study were generated through either experimental in-vitro measurements, involving the recording of neuronal electrical activity using a 60-channel MEA, or through in-silico simulations using the NEST simulator. Both datasets underwent a similar preprocessing approach, where only significant culture activity was incorporated into the training dataset. This involved the application of burst and NB detection algorithms [31, 32]. The experimental dataset required additional preprocessing before burst detection, entailing the transformation of raw electrical signals into spike trains using filtering and a spike detection algorithm [33]. A comprehensive description of the experimental procedures and the algorithms employed in this study can be accessed in the Supplementary Materials section. Additionally, details regarding the in-silico simulation are provided in Section 4.6.
@@ 4.3.1 Circuit Integration Time and Data Binning
A fundamental principle underpinning the RC model is the utilization of rate coding for the transmission of neural information. Consequently, the training data was structured as a time sequence representing the network state, encoded in spike-rate values. This approach was based on the assumption that there exists a characteristic range of neural spike integration time (found to be within a timescale of a few milliseconds), defining the information processing duration for individual circuits or populations within the system. We empirically demonstrated a close correspondence between the integration time (yielding optimal performance of the RC model) and the peak of the inter-burst-interval (IBI) histogram within the 1-10 millisecond range. This IBI histogram was constructed using intervals between consecutive bursts across all channels, employing the procedure described in the Supplementary Materials, also used for the inter-spike-interval (ISI) histogram. The spike traces are subsequently segmented into time bins of width corresponding to integration time, where each bin contains the ISR. Accordingly, we transform the spike trains data to ISR data:
@@ 4.3.2 Training and Validation Datasets
Following the steps described in the previous subsection, we generate the training and validation datasets. To ensure a consistent domain for the RC model across all datasets, we standardize the ISR data within the $\mathbf{\Sigma}$ matrix, scaling its values to fall within the range of 0 to 1. Subsequently, we construct the dataset by extracting data batches from $\mathbf{\Sigma}$ in segments from the full-length data, employing the following approach:
$\overline{\mathbf{\Sigma}}\Bigl{[}i,n_{j}:n_{j}+N_{j}+\Delta-1\Bigr{]}\rightarrow\mathbf{Y}_{j}\left[i,1:L_{NB}^{j}+\Delta\right]$ (18)
$\overline{\mathbf{\Sigma}}$ is the normalized $\mathbf{\Sigma}$ matrix; $n_{j}$ refers to the time-bin which indicates the beginning of the $j$th NB. $N_{j}$ denotes the duration of the $j$th NB, specified in the number of time bins. The parameter $\Delta$ is used to accommodate extra time steps, guaranteeing the inclusion of the latter part of the NB. Accordingly, the matrix $\mathbf{Y}_{j}$ represents the training batch number $j$.
@@
All batches $\mathbf{Y}^{tr}_{j}$ are then provided to the lasso regression algorithm (Eq. (2.1)) as merged sequence $\mathbf{Y}^{tr}$ (merging all matrices $\mathbf{Y}^{tr}_{j}$ along time dimension). The training however is informed about a beginning of a new batch, by resetting the time step to $n=0$ and the resrevoir state to $\mathbf{x}[0]=0$ (Eq. (2b)); in such way we decouple effects of one NB on another, taking in consideration that the time interval between different NBs is considerably longer than the interaction between the populations. The remaining validation batches $\mathbf{Y}^{val}$ are then used to evaluate the training, by calculating the validation loss:
@@
$y_{i}^{val},\tilde{y}_{i}^{val}$ are the ground truth and modeled validation data for each channel $i$; $\bar{W}_{out},\bar{b}$ (Eq. (19c)) are the trained weights of the output matrix $\mathcal{W}_{out}$ and bias vector $\mathbf{b}$ (Eq. (2.1)); To enhance the robustness of our predictions, we address the issue arising from a significant portion of zero sequences in the traces $y,\tilde{y}$. The prevalence of these zero sequences can potentially lead to overly simplistic predictions. To mitigate this, we incorporate weights $\omega$ (Eq. (19b)) into our loss estimations. These weights, defined in Eq. (19d), are designed to assign greater importance to non-zero predictions within the trace, thereby addressing the potential bias introduced by the prevalence of zero sequences and enhancing the sensitivity of the model to more informative elements in the data.

### 05 Mar 2024 v3

@@ 5.1 The paradigm
We consider a multi-site measurement of electrophysiological signals from a neuronal culture, such as 2D microelectrode array (MEA). We seek to represent the tested culture as a network where each node corresponds to one measurement electrode. Each electrode samples the electrophysiological signals from the neuron ensemble (consisting of a few neurons) found in its vicinity (Fig. 0(a)). Therefore, each node has to represent a complex neuronal circuit whose dynamics by itself is driven by numerous interacting neurons. We hence define the domain of the MEA measurement as the macroscopic domain, which is the one described by the network. The neuronal structure probed by each electrode will be referred as the microscopic domain (or later as the reservoir domain) and corresponds to the single node in the network. The data unit which is contained in each of these nodes is a sample of the electrophysiological signals measured by one electrode expressed in instantaneous spike-rate measured in a specified time window. The time window is determined by a characteristic information flow rate in the culture. This unit of time is dependent on many properties of the culture such as neuron density in the culture, age of the culture and other [31], and it characterizes the signal integration time of each electrode.
@@
Let us represent the macro-domain state of the network at each time step $n=1,2,3...$ with a vector $\mathbf{y}[n]$, where each component of the vector describes the state of a single node, i.e., $\mathbf{y}[n]$ is the signal representation of each electrode at time $n$. Our aim is to seek the time propagation operator (or a synaptic transmission function) $\hat{\mathcal{F}}$ (Eq. (3)) transforming state $\mathbf{y}[n]$ to $\mathbf{y}[n+1]$. The operator $\hat{\mathcal{F}}$, which is likely to be non-linear due to the nature of neuronal networks, should describe as closely as possible the experimental observation in the electrophysiological measurements, i.e., we aim to fit a model to an observation which can model or predict the spatio-temporal patterns of the neuronal activity in the culture under test.
@@
We then consider that each node in the macro-domain network represents a complex neuronal signal-processing unit. It arises from the fact that typically every measurement site is surrounded by neurons which may be as many as a dozen. The morphology and functionality of each of these micro-circuits embedded in each node of the macro-domain network cannot be easily obtained from the electrophysiological measurements. Also modeling of such neuronal structures is not as easy and has been studied for decades, with numerous models for different scales of dimensions and time [4, 5].
@@
Hence our approach is to represent each measurement node (electrode) as a gate to a particular neuronal circuit (reservoir), where the signals measured at each node are an outcome of a complex operation involving each circuit and the whole network. We therefore propose the artificial neural network (ANN) structure depicted in Fig. 0(c). This structure represents a simplification of the neuronal dynamics, where each of the neuronal circuits is a black box, whose morphology and functionality are not known but assumed to be reasonably random.
@@
As seen in Fig. 0(c) and Figure Supplementary 1, we assume that the signal sampled at each node is an input to and an output from a higher dimensional domain with specific connectivity and functionality. Each reservoir, associated with a node in the macro-domain, represents a micro-neuronal circuit embedded at each of the measurement sites and has inner interconnections that represent the connectivity of the micro-circuits. Each such circuit performs a nonlinear transformation, creating an updated reservoir state, which on one side is stored as a memory to be integrated into the next time steps, and on the other side is used to form the next state of the macro-domain network.
@@ 5.2.1 Domains and Dimensions
We denote by $N_{ch}$ the number of nodes in the network, where each node is directly associated with an electrode (or a channel) in the experimental measurement. Assuming that the neurons are uniformly distributed in the culture, we appoint a fixed number of connections between each node and the corresponding micro-circuit, such that for each node of the network there is one micro-reservoir (see Fig. 0(c)). Defining $m$ the dimension of the micro-reservoir, it results that the dimension of the reservoir layer $N_{res}$ is:
$N_{res}=m\times N_{ch}.$ (10)
It follows that each $m$ components in the vector space of the reservoir layer correspond to one node in the macro domain. We may associate $m$ with the relative size of each micro-circuit.
@@ 5.2.2 Input Layer
The input layer refers to the stage between the macro domain and the reservoir layer. Here we assume that the input state to the reservoir is a linear transformation of the data at the corresponding node, such that each component in the macro-domain transforms directly to $m$ inputs of $N_{ch}$ micro-reservoirs in the reservoir layer, and described the inner dynamics of a single micro-circuit. This is done with the linear transformation described by Eq. (2a), where the state $\mathbf{y}\in\mathbb{R}^{N_{ch}\times 1}$ is transformed to the input reservoir state $\mathbf{x}_{in}\in\mathbb{R}^{N_{res}\times 1}$ through an input matrix $\mathcal{W}_{in}\in\mathbb{R}^{N_{res}\times N_{ch}}$. Since $\mathcal{W}_{in}$ maps each node to a corresponding micro-reservoir, it is represented by the following matrix:
${\mathcal{W}_{in}=\begin{bmatrix}\left(\mathbf{w}_{in}^{(1)}\right)&0&0&\cdots&0\\ 0&\left(\mathbf{w}_{in}^{(2)}\right)&0&\cdots&0\\ 0&0&\left(\mathbf{w}_{in}^{(3)}\right)&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\\ 0&0&0&&\left(\mathbf{w}_{in}^{{(N_{ch})}}\right)\end{bmatrix}}$ (11)
@@ 5.2.3 Reservoir Layer
The reservoir layer contains $N_{ch}$ independent micro-circuits (reservoirs) with $m$ nodes each. Each such reservoir models the neuronal circuit probed by each electrode. This layer has two main functionalities: i) nonlinear time-operator, and ii) reservoir state integrator. The dynamics of the reservoir layer is recurrent. At each time step $n$ the reservoir state $\mathbf{x}[n-1]$ is nonlinearly mapped into a new reservoir state $\mathbf{x}[n]$ (Eq. (2b)). This discrete differential relation provides cumulative data at each time step and carries the temporal memory on the activity of the network. Note that at time step $n$, the reservoir state $\mathbf{x}[n-1]$ undergoes an inner transformation through the reservoir matrix $\mathcal{W}_{res}\in\mathbb{R}^{N_{res}\times N_{res}}$, then integrated with the reservoir input state $\mathbf{x}_{in}$, followed by a nonlinear operation $\mathbf{f}_{NL}$. The operation of $\mathcal{W}_{res}$ is linear and describes the inner data propagation between the nodes of the reservoir layer. The synaptic morphology of each micro-reservoir is random and since they are independent, there are no coupling elements between them. In addition, we assume that $\mathcal{W}_{res}$ preserves the energy of the state, hence it is represented by a block-diagonal and orthogonal matrix as the following:
@@
Furthermore, we note in Eq. (2b) the memory parameter $10<\alpha<$ 1, which factorizes the orthogonal matrix $\mathcal{W}_{res}$. This represents the memory strength, such that the effect of the reservoir state $\mathbf{x}[n]$ on the state $\mathbf{x}[n+k]$ will scale as $\alpha^{k}$. $\alpha=0$ indicates that the system is memoryless and the current state at time step $n$ depends only on the input.
The nonlinear function $\mathbf{f}_{NL}$ in Eq. (2b) accounts for the impact of the network saturation and plasticity. Specifically, when subjected to a persistent input, the reservoir tends to saturate, while in the absence of input, it retains an excited state due to the influence of previous signals (plasticity). Note that the plasticity persistence effect depends on the parameter $\alpha$. We tried different nonlinear functions, such as tanh and sigmoid, but the performances of the model showed a weak dependence on the choice of the function. For the results presented in this paper, we considered:
@@
where $\Theta(\mathbf{\xi})$ is the unit step function and is used to limit the emerging values from the reservoir layer to positive values only. Furthermore, we introduced in Eq. (2b) the factor $\hat{\mathcal{S}}$ to give each micro-reservoir neuron a different nonlinear response strength. $\hat{\mathcal{S}}$ does not couple different neurons, $\hat{\mathcal{S}}\in\mathbb{R}^{N_{res}\times N_{res}}$ and is a diagonal matrix containing random normally-distributed synaptic strengths on its diagonal.
@@ 5.2.4 Output Layer
The output layer transforms the reservoir state back to the macroscopic domain and is described by Eq. (2c). Here, we assume a fully connected layer, such that all the $N_{res}$ reservoir nodes are weighted and connected to all the $N_{ch}$ nodes of the macroscopic network. This layer expresses the synaptic connectivity between the different nodes of the network. It is assumed that this transformation is purely linear, taking into consideration that the overall nonlinearity of the model is dominated by the reservoir layer. Unlike $\mathcal{W}_{in}$, $\mathcal{W}_{res}$ and $\hat{\mathcal{S}}$, which are matrices with random and constrained elements, $\mathcal{W}_{out}$ has no constraints on the values of its elements, rather its values are learned during the training by using a Lasso regression (Eq. (2)).
@@ 5.3 Data
The model operates on the principles of a rate-coded spiking neural network. For model training, we generated tailored datasets that govern episodes of network bursting (NB) activity, representing them as multidimensional sequences of ISR data. The data in this study were generated through either experimental in-vitro measurements, involving the recording of neuronal electrical activity using a 60-channel MEA, or through in-silico simulations using the NEST simulator. Both datasets underwent a similar preprocessing, where only significant activities were incorporated into the training dataset. This involved the application of burst and NB detection algorithms [36, 37]. The experimental data required additional preprocessing before burst detection, entailing the transformation of raw electrical signals into spike trains using filtering and a spike detection algorithm [38]. A comprehensive description of the experimental procedures and the algorithms employed are given in the Supplementary Materials.
@@ 5.3.1 Circuit Integration Time and Data Binning
The training data was structured as a time sequence representing the network state, encoded in spike-rate values. This approach was based on the assumption that there exists a characteristic range of neural spike integration time (found to be within a timescale of a few milliseconds), defining the information processing duration for individual populations within the neuronal culture (Fig. 0(b)). We empirically demonstrated a close correspondence between the integration time (yielding optimal performance of the RC model) and the peak of the inter-burst-interval (IBI) histogram (Fig. 4(d)) within the 2-10 millisecond range. The spike traces are subsequently segmented into time bins of duration equal to the integration time, where each bin contains the ISR (see details in Sec. II.b of the Supplementary Materials). Accordingly, we transform the spike trains data to ISR data:
@@ 5.3.2 Training and Validation Datasets
Following the steps described in the previous subsection, we generate the training and validation datasets. To ensure a consistent domain for the RC model across all datasets, we standardize the ISR data within the $\mathbf{\Sigma}$ matrix, scaling its values to fall within the 0 to 1 range. Subsequently, we construct the dataset by extracting data batches from $\mathbf{\Sigma}$ in segments from the full-length data, employing the following approach:
$\overline{\mathbf{\Sigma}}\Bigl{[}i,n_{j}:n_{j}+N_{j}+\Delta-1\Bigr{]}\rightarrow\mathbf{Y}_{j}\left[i,1:N_{j}+\Delta\right]$ (18)
$\overline{\mathbf{\Sigma}}$ is the normalized $\mathbf{\Sigma}$ matrix; $n_{j}$ refers to the time-bin which indicates the beginning of the $j$-th NB. $N_{j}$ denotes the duration of the $j$-th NB, specified in the number of time bins. The parameter $\Delta$ is used to accommodate extra time steps, guaranteeing the inclusion of the latter part of the NB. Accordingly, the matrix $\mathbf{Y}_{j}$ represents the training batch number $j$.
@@
All batches $\mathbf{Y}^{tr}_{j}$ are then provided to the Lasso regression algorithm (Eq. (2)) as merged sequence $\mathbf{Y}^{tr}$ (merging all matrices $\mathbf{Y}^{tr}_{j}$ along time dimension). The training however is informed about the beginning of a new batch, by resetting the time step to $n=0$ and the reservoir state to $\mathbf{x}[0]=0$ (Eq. (2b)). In this way, we decouple the effects of one NB on another since the time interval between different NBs is considerably longer than the neuronal interaction time. The remaining validation batches $\mathbf{Y}^{val}$ are then used to evaluate the training, by calculating the validation loss:
@@
$y_{i}^{val},\tilde{y}_{i}^{val}$ are the observed and modeled validation data for each node $i$; $\bar{W}_{out},\bar{b}$ (Eq. (19c)) are the learned elements of $\mathcal{W}_{out}$ (i. e. the trained weights) and of the bias vector $\mathbf{b}$ (Eq. (2)). To enhance the robustness of our predictions, we use $\omega$ (Eq. (19b)) in our loss estimations. They mitigate the possible presence of a significant portion of zero sequences in the traces $y,\tilde{y}$. These zero sequences can potentially lead to overly simplistic predictions. These $\omega$ weights, defined in Eq. (19d), assign greater importance to non-zero predictions within the trace, thereby addressing the potential bias introduced by the prevalence of zero sequences and enhancing the sensitivity of the model to more informative elements in the data.